OpenStudy (anonymous):

Solve the inequality 5/8 < (3/8)x - 1/4 ?

OpenStudy (bahrom7893):

5<3x-2 7/3<x x>7/3

OpenStudy (anonymous):

(3/8)x > 5/8 + 1/4 = 7/8 x > 7/8 * 8/3 = 7/3

OpenStudy (anonymous):

x > 7/3

OpenStudy (anonymous):

fractions, ick. multiply by 8 and get \[5<3x-2\] \[7<3x\] \[\frac{7}{3}<x\]

OpenStudy (anonymous):

i've noticed you dont like fractions satellite! lol

OpenStudy (bahrom7893):

that is exactly what i did..

OpenStudy (bahrom7893):

lol